Four balls are to be randomly chosen from an urn containing 4 red, 5 green, and 6 blue balls.

1. Find the probability that at least one red ball is chosen?

2. Given that no red balls are chosen, what is the probability that there are exactly 2 green balls among the four balls chosen.
Answer #1

Red balls =4 Green balls = 5 Blue balls = 6 ls = 4+5+ b = 15 4 4 balls are randomly selected. B At least I red ball) = 1- P (